Using the Python to do int operators +, -, *, //, and ** (not %), and the numbers 2, 3, 4, and 5, and an expression

that uses all four numbers exactly once, uses any three of the operators exactly once, and evaluates to 26.
Answer #1

a=2

b=3

c=4

d=5

result=d**a-b+c

# first d**a is evaluated and it is having highest precedence 5**2=25 and +,- are having same prioirity

so its associativity is from left to right. Therefore 25-b=25-3=22+c=22+4=26

print(result)